OpenStudy (anonymous):

Which ratio is proportional to 4/15 ? a. 12/45 b. 2/7 c. 16/30 d. 36/75

OpenStudy (anonymous):

A

OpenStudy (anonymous):

a) 12/45=(12/3)/(45/3)=4/15

OpenStudy (anonymous):

The reason it is A is because: 12/45=2.66666.... and 4/15=2.6666666.... They equal the same thing

OpenStudy (anonymous):

A, 4/15 multiplied with 3 equals 12/45